Hi can anyone out there please help me with this word problem.
A piece of rope 5 m long is cut into two pieces so that one piece is 3/5ths as long as the other. Find the length of each piece.

Answer by edjones(7569)

3x/5 + x = 5
3x+5x=25 multiply each side by 5
8x=25
.
x=25/8 m
3x/5=15/8 m
.
Ed